### v2

@@ A.8 Proof of Proposition 5
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=s\right.\right]$
$=\delta_{j,k}(s)~{}\mathbb{E}\left[g(Y)\left|Q_{j}=(j,k),\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=s\right.\right]$
$\hskip 14.22636pt+(1-\delta_{j,k}(s))~{}\mathbb{E}\left[g(Y)\left|Q_{j}\neq(j,k),\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=s\right.\right]$
$=\delta_{j,k}(s)~{}\mathbb{E}\left[g(Y(j))\left|Q_{j}=(j,k),S_{j}=s\right.\right]$
$\hskip 14.22636pt+(1-\delta_{j,k}(s))~{}\mathbb{E}\left[g(Y)\left|Q_{j}\neq(j,k),\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=s\right.\right],$
@@
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{+}\right.\right]$
$=\delta_{j,k}^{+}~{}\mathbb{E}\left[g(Y(j))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$
$\hskip 14.22636pt+(1-\delta_{j,k}^{+})~{}\mathbb{E}\left[g(Y)\left|Q_{j}\neq(j,k),\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{+}\right.\right],$
@@
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{-}\right.\right]$
$=\delta_{j,k}^{-}~{}\mathbb{E}\left[g(Y(k))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$
$\hskip 14.22636pt+(1-\delta_{j,k}^{-})~{}\mathbb{E}\left[g(Y)\left|Q_{j}\neq(j,k),\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{-}\right.\right].$
We know the expectations on the left-hand sides of the last two equations, but we do not know the $\delta$s or the expectations on the right-hand sides. Above the cutoff, the goal is to partially identify $\mathbb{E}\left[g(Y(j))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ using the distribution of $g(Y)$ conditional on $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$ plus knowledge of a strictly positive lower bound on $\delta_{j,k}^{+}$, i.e., $\underline{\delta}_{j,k}^{+}$. Likewise, below the cutoff, the goal is to partially identify $\mathbb{E}\left[g(Y(k))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ using the distribution of $g(Y)$ conditional on $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{-}$ plus $\underline{\delta}_{j,k}^{-}$. This problem fits the setting of horowitz1995, specifically, Propositions 1–4 and Corollary 4.1 of their paper.
Part (i)
Assume $g(Y)$ is a continuous random variable. Assume for the time being that we know the fraction of individuals with $Q_{j}=(j,k)$, that is, $\delta_{j,k}^{+}>0$. The lowest possible value for the mean $\mathbb{E}[g(Y(j))|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs when all individuals with $Q_{j}=(j,k)$ are in the lower tail of the distribution of $g(Y)$ conditional on $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$. This gives the lower bound
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)<F_{j,k+}^{-1}(\delta_{j,k}^{+}),S_{j}=c_{j}^{+}\right.\right].$
On the other hand, the highest possible value for the mean $\mathbb{E}\left[g(Y(j))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ occurs when all individuals with $Q_{j}=(j,k)$ are in the upper tail of the distribution of $g(Y)$ conditional on $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$. This gives the upper bound
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)>F_{j,k+}^{-1}(1-\delta_{j,k}^{+}),S_{j}=c_{j}^{+}\right.\right].$
We do not know $\delta_{j,k}^{+}$, but we know that $0<\underline{\delta}_{j,k}^{+}\leq\delta_{j,k}^{+}$, so all values $\delta_{j,k}^{+}\in[\underline{\delta}_{j,k}^{+},1]$ are possible. As the fraction $\delta_{j,k}^{+}$ decreases, the lower bound derived above decreases, and the upper bound increases. Thus, the widest bounds occur when $\delta_{j,k}^{+}=\underline{\delta}_{j,k}^{+}$, that is,
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)<F_{j,k+}^{-1}(\underline{\delta}_{j,k}^{+}),S_{j}=c_{j}^{+}\right.\right]$
@@
$\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)>F_{j,k+}^{-1}(1-\underline{\delta}_{j,k}^{+}),S_{j}=c_{j}^{+}\right.\right].$
The bounds for $\mathbb{E}\left[g(Y(k))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ are derived in an analogous fashion.
Part (ii)
Assume $g(Y)=Y$ and $Y$ is binary. The expressions for the bounds in Part (i) are not valid in this case. To see this, take the lower bound for the mean of $Y(j)$ as an example. The quantile function $F_{j,k+}^{-1}(\underline{\delta}_{j,k}^{+})$ equals either 00 or $1$, so that the conditioning event $Y<F_{j,k+}^{-1}(\underline{\delta}_{j,k}^{+})$ is either empty or has probability that is generally different from $\underline{\delta}_{j,k}^{+}$.
To start, note that
$\mathbb{E}\left[g(Y(j))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]=\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right].$
Define
$q_{j,k}^{+}:=\mathbb{E}\left[Y\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{+}\right.\right]=\mathbb{P}\left[Y=1\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{+}\right.\right].$
Assume for the time being that we know $\delta_{j,k}^{+}$.
Case (a): Suppose $1-q_{j,k}^{+}\geq\delta_{j,k}^{+}$, that is, that the fraction of individuals with $Y=0$ in the population of individuals with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$ is greater than or equal to the fraction of individuals with $Q_{j}=(j,k)$ in that same population. In this case, the lowest possible value for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ occurs when all $\delta_{j,k}^{+}$ individuals have $Y$ equal to 00. That lowest value is zero.
Case (b): Now, suppose $1-q_{j,k}^{+}<{\delta}_{j,k}^{+}$. It is no longer possible to have all ${\delta}_{j,k}^{+}$ individuals with $Y$ equal to 00. The lowest possible value for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ has as many of ${\delta}_{j,k}^{+}$ individuals as we can with $Y=0$, that is, $1-q_{j,k}^{+}$ of them. The remaining ${\delta}_{j,k}^{+}-(1-q_{j,k}^{+})$ individuals must have $Y=1$. That lowest possible value is $(1-q_{j,k}^{+})/{\delta}_{j,k}^{+}\times 0~{}+~{}[{\delta}_{j,k}^{+}-(1-q_{j,k}^{+})]/{\delta}_{j,k}^{+}\times 1=1-(1-q_{j,k}^{+})/\delta^{+}_{j,k}$.
The lower-bound expression for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ that covers both cases (a) and (b) is as follows:
$\max\left\{1-\frac{1-q_{j,k}^{+}}{{\delta}^{+}_{j,k}}~{},~{}0\right\}.$
Next, we consider the upper bound for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$.
Case (c): Suppose $q_{j,k}^{+}\geq{\delta}_{j,k}^{+}$, that is, that the fraction of individuals with $Y=1$ in the population of individuals with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$ is greater than or equal to the fraction of individuals with $Q_{j}=(j,k)$ in that same population. The highest possible value for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ occurs when all ${\delta}_{j,k}^{+}$ individuals have $Y=1$. That highest value is $1$.
Case (d): Now, suppose $q_{j,k}^{+}<{\delta}_{j,k}^{+}$. It is no longer possible to have all ${\delta}_{j,k}^{+}$ individuals with $Y$ equal to $1$. The highest possible value for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ has as many of the ${\delta}_{j,k}^{+}$ individuals as we can with $Y=1$, that is, $q_{j,k}^{+}$ of them. The remaining ${\delta}_{j,k}^{+}-q_{j,k}^{+}$ individuals must have $Y=0$. That highest possible value is $q_{j,k}^{+}/{\delta}_{j,k}^{+}\times 1~{}+~{}({\delta}_{j,k}^{+}-q_{j,k}^{+})/{\delta}_{j,k}^{+}\times 0=q_{j,k}^{+}/\delta^{+}_{j,k}$.
The upper-bound expression for $\mathbb{P}\left[Y(j)=1\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]$ that covers both cases (c) and (d) is as follows:
$\min\left\{\frac{q_{j,k}^{+}}{{\delta}^{+}_{j,k}}~{},~{}1\right\}.$
We do not know $\delta_{j,k}^{+}$; we know only the lower bound $\underline{\delta}_{j,k}^{+}>0$. As in Part (i) above, we compute the bounds at $\underline{\delta}_{j,k}^{+}$ because they grow weakly wider as $\delta_{j,k}^{+}$ decreases.
The bounds for $\mathbb{P}[Y(k)=1|Q_{j}=(j,k),S_{j}=c_{j}]$ are derived in an analogous fashion.
@@
Appendix B Sharp Bounds on Treatment Effects
In this section, we utilize Artstein’s inequality (Theorem A.1 in molinari2020) to characterize sharp bounds on the joint distribution of potential outcomes and true local preferences. That set of distributions produces sharp bounds on the averages of the treatment effects $Y(j)-Y(k)$ conditional on $Q_{j}=(j,k)$ and $S_{j}=c_{j}$ for any comparable pair $(j,k)\in\mathcal{P}$. Recall that we denote the space of possible outcomes $Y$ as $\mathcal{Y}$. Let $2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}$ denote the power set of ${\mathcal{J}^{0}\times\mathcal{J}^{0}}$ and $2^{\mathcal{Y}}$ denote the power set of ${\mathcal{Y}}$.
Theorem 1.
Proof of Theorem 1: The random set ($(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a measurable map from $\Omega$ to $\mathcal{Y}\times\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 from molinari2020, we say that ($(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a random closed set because, for every compact set $\mathcal{K}\in\mathbb{R}^{3}$, the set $\{\omega\in\Omega:(\{Y(\omega)\}\times\bm{\mathrm{Q}}_{j}(\omega))\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $(Y,Q_{j})$ and the random set ($(\{Y\}\times\bm{\mathrm{Q}}_{j})$ are measurable maps on the same probability space, and $\mathbb{P}\left[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 in molinari2020) characterizes the sharp set of all possible probability distributions for $(Y,Q_{j})$ that are consistent with our observation of ($(\{Y\}\times\bm{\mathrm{Q}}_{j})$ and the fact that $\mathbb{P}[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s]=1$. For any $s\in\mathcal{S}_{j}$, the inequality says that
$\mathbb{P}\left[Y\in A,\bm{\mathrm{Q}}_{j}\subseteq B|S_{j}=s\right]\leq\mathbb{P}\left[Y\in A,Q_{j}\in B|S_{j}=s\right]\;\;\;\forall A\in 2^{\mathcal{Y}},\;\ B\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}.$ (26)
Next, we use the cutoff characterization and rewrite the right-hand side of (26) as a sum. For $s\geq c_{j}$,
$=\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b)\in A,Q_{j}=(b,b^{\prime})|S_{j}=s\right].$ (27)
Substitute (27) into (26) and take the limit as $s\downarrow c_{j}$ on both sides,
$\mathbb{P}\left[Y\in A,\bm{\mathrm{Q}}_{j}\subseteq B|S_{j}=c_{j}^{+}\right]\leq\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b)\in A,Q_{j}=(b,b^{\prime})|S_{j}=c_{j}\right]\;\;\;\forall A\in 2^{\mathcal{Y}},\;\ B\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}},$ (28)
Similarly, for $s>c_{j}$,
$=\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b^{\prime})\in A,Q_{j}=(b,b^{\prime})|S_{j}=s\right].$ (29)
Use (29) into (26), and take the limit as $s\uparrow c_{j}$ on both sides,
$\mathbb{P}\left[Y\in A,\bm{\mathrm{Q}}_{j}\subseteq B|S_{j}=c_{j}^{-}\right]\leq\sum_{(b,b^{\prime})\in B}\mathbb{P}\left[Y(b^{\prime})\in A,Q_{j}=(b,b^{\prime})|S_{j}=c_{j}\right]\;\;\;\forall A\in 2^{\mathcal{Y}},\;\ B\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}.$ (30)
Lemma 1 of chesher_ecta2017 applied to our case shows that inequalities (28) and (30) are equivalent to:
$\square$
Theorem 1 characterizes the sharp set of all possible probability values of $\mathbb{P}\left[Y(d)\in A,Q_{j}=(b,b^{\prime})|S_{j}=c_{j}\right]$ for $A\subseteq\mathcal{Y}$, $(b,b^{\prime})\in\mathcal{J}^{0}\times\mathcal{J}^{0}$, and $d\in\{b,b^{\prime}\}$. For a fixed $g\in\mathcal{G}$ of Assumption 2, that set of distributions allows us to define the sharp set of all possible means of potential outcomes,
$\mathbb{E}\left[g(Y(d))|Q_{j}=(j,k),S_{j}=c_{j}\right],$
$\mathbb{E}\left[g(Y(j))-g(Y(k))|Q_{j}=(j,k),S_{j}=c_{j}\right].$
In case of continuous $Y$, it is impossible to directly evaluate all inequalities of Theorem 1 because there are uncountably many sets $A\in 2^{\mathcal{Y}}$. This is one of the drawbacks of the Artstein inequality approach, which has been extensively discussed by beresteanu_joe2012. In case $Y$ takes finitely many values, e.g., when $Y$ is binary, the number of such inequalities is feasible to evaluate because $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup+}_{j}$ (or $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup-}_{j}$) has finitely many elements. In fact, the number of inequalities is slightly higher than the number of inequalities in Proposition 4 of the main text, which we utilize to compute lower bounds on $\delta_{j,k}^{+}$ and $\delta_{j,k}^{-}$.
@@ Appendix C Empirical Appendix
Table 3: Descriptive statistics: Students
All students Applicants to medicine
at PUC Santiago
Mean Std. dev. Mean Std. dev.
Number of programs in ROL 4.86 2.20 5.18 2.01
ROL strictly shorter than permitted .80 .40 .77 .42
Assigned (to any prog.) .65 .48 .68 .46
Rank of assigned program, cond. on assigned 2.24 1.59 2.58 1.68
Reapplies .19 .39 .25 .43
Graduates from assigned prog., cond. on assigned .22 .41 .32 .48
Graduates from any program .77 .42 .89 .31
Number of students 519,409 9,398
The left panel of the table provides descriptive statistics on the population of participants in the centralized college assignment mechanism in Chile between 2004 and 2010. The right panel provides descriptive statistics on the subpopulation of students who included medicine at PUC Santiago (our program $j$ of interest in Section 5.2) in their rank-ordered list (ROL).
Table 4: Descriptive statistics: Programs
All programs Medicine
at PUC Santiago
Mean Std. dev. Mean Std. dev.
Number of applicants (per year) 474 351 1,094 138
Number of admitted students (per year) 66 48 90 .90
Year-to-year absolute change in cutoff (points) .85 6.9 2.5 11.5
Cutoff 720 10
@@
The left panel of the table provides descriptive statistics on all programs involved in the centralized college assignment mechanism in Chile between 2004 and 2010. The right panel provides descriptive statistics on the medicine program at PUC Santiago (our program $j$ of interest in Section 5.2).

### v3

@@ B.7 Proof of Proposition 5
[…]
@@
[…]
@@
[…]
Binary Outcome.
[…]
@@
[…]
Continuous Outcome.
[…]
and
[…]
@@ Appendix C Empirical Appendix
Descriptive statistics.
Tables C-1 and C-2 provide descriptive statistics about the programs used as case studies in Section 5.2 and their applicants.
Table C-1: Descriptive statistics: Programs
All programs Medicine Bachillerato
at PUC at UChile
Mean S.dev. Mean S.dev. Mean S.dev.
Total applicants (per year) 474 351 1,094 138 2430 317
Total admitted stud. (per year) 66 48 90 .90 385 39
Year-to-year absolute change in cutoff .85 18.2 2.5 11.5 2.0 6.8
Cutoff 780 7.4 670 4.0
@@
The first column of the table provides descriptive statistics on all programs involved in the centralized college assignment mechanism in Chile between 2004 and 2010. The second and third columns provide descriptive statistics on the medicine program at PUC Santiago and the Bachillerato de Ingreso Comun at UChile (two of our programs of interest in Section 5.2).
Table C-2: Descriptive statistics: Students
All students Medicine Bachillerato
at PUC at UChile
Mean S.dev. Mean S.dev. Mean S.dev.
Number of programs in ROL 4.86 2.20 5.18 2.01 5.33 1.84
ROL strictly shorter than permitted .80 .40 .77 .42 .79 .41
Assigned (to any prog.) .65 .48 .68 .46 .71 .45
Rank of assigned prog., cond. on assigned 2.24 1.59 2.58 1.68 2.64 1.59
Reapplies .19 .39 .25 .43 .31 .46
Graduates from any program .77 .42 .90 .31 .83 .37
Grad. from assigned prog., cond. assigned .22 .41 .32 .48 .16 .37
Number of students 519,409 6,438 14,571
The first column of the table provides descriptive statistics on the population of participants to the centralized college assignment mechanism in Chile between 2004 and 2010. The second and third columns provide descriptive statistics on the subpopulation of students who included, respectively, medicine at PUC Santiago and Bachillerato de Ingreso Comun at UChile (two of our programs of interest in Section 5.2) in their rank-ordered list (ROL).
Average structural functions and the importance of preferences for graduation outcomes: comparing applicants across cutoffs in Figure 3.
[…]
Table C-3: Relative average distance between students across cutoffs $c_{j}$
$j_{2}$
Medicine Medicine Odontology Kinesiology Math
UChile U.Sant. de Ch. UChile UChile PUC
Medicine, UChile 1 .98 .98 1 1.13
Med., U.Sant. de Ch. 1.09 1 1.02 .98 1.13
$j_{1}$ Odontology, UChile 1.05 .98 1 .97 1.11
Kinesiology, UChile 1.16 1.03 1.06 1 1.13
Math, PUC 1.1 .99 1.01 .95 1
[…]
Table C-4: Local samples and estimated $\underline{\delta}_{j,k}$s for $k=$ Bachillerato de Ingreso Comun at UChile
(1) (2) (3) (4)
SPO WPO WPO WPO
UMAS UMAS
fields
$c_{j}$ $\geq c_{j}$ $c_{j}$ $\geq c_{j}$ $c_{j}$ $\geq c_{j}$ $c_{j}$ $\geq c_{j}$
Local first-best $j$: Medicine at U.Chile
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 155 108 140 181 1001 370 1007 736
Share listing <<8 choices .38 .38 .8 .96 .9 .97 .9 .91
$\underline{\delta}_{j,k}$ 1 .43 .92 .83 .89 .28 .89 .04
Local first-best $j$: Medicine at U.Santiago
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 104 77 201 34 239 89 241 334
Share listing <<8 choices .75 .18 .89 .91 .89 .94 .89 .81
$\underline{\delta}_{j,k}$ .84 .58 .43 1 .36 .51 .36 .13
Local first-best $j$: Odontology at U.Chile
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 127 63 111 111 936 250 957 316
Share listing <<8 choices .14 .41 .72 .86 .89 .9 .88 .88
$\underline{\delta}_{j,k}$ .71 .42 .23 .68 .18 .29 .16 .07
Local first-best $j$: Kinesiology at U.Chile
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 118 28 307 30 544 94 622 151
Share listing <<8 choices .33 .53 .87 .83 .89 .93 .87 .91
$\underline{\delta}_{j,k}$ .39 1 .15 1 .08 .34 .07 .2
Local first-best $j$: Math at PUC
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 50 86 36 235 36 235 148 352
Share listing <<8 choices .51 .68 .72 .92 .72 .92 .83 .92
$\underline{\delta}_{j,k}$ 1 .71 1 .15 1 .15 .24 .09
The table describes the sample used in the estimation of the effect on outcomes of interest of assignment to Bachillerato de Ingreso Comun at UChile ($k$ in the table), under different sets of assumptions used to construct $\bm{\mathrm{Q}}_{j}$. Given each set of assumptions (shown at the top of each column), the table shows the number of students within a 30-point bandwidth on either side of the cutoff to program $j$ for whom $\bm{\mathrm{Q}}_{j}$ contains the pair $(j,k)$ and the share of these students who listed strictly fewer than eight choices in their ROLs. For each pair $(j,k)$ and each set of assumptions, the table also shows the estimated $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ used in the estimation of the bounds on the treatment effects. In the derivation of the bounds for the average structural function shown in Figure 3, only $\underline{\delta}_{j,k}^{-}$ is used, as shown in Eq. (B.7). We still include $\underline{\delta}_{j,k}^{+}$ for completeness and illustration purposes as the same logic applies to both $\underline{\delta}_{j,k}^{-}$ and $\underline{\delta}_{j,k}^{+}$.
Table C-5: Local samples and estimated $\underline{\delta}_{j,k}$s for $j=$ medicine at PUC
(1) (2) (3) (4)
SPO WPO WPO WPO
UMAS UMAS
fields
$c_{j}$ $\geq c_{j}$ $c_{j}$ $\geq c_{j}$ $c_{j}$ $\geq c_{j}$ $c_{j}$ $\geq c_{j}$
Next-preferred $k$: medicine at UChile
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 382 390 537 427 537 427 537 427
Share listing <<8 choices .88 .92 .91 .93 .91 .93 .91 .93
$\underline{\delta}_{j,k}$ 1 .92 .67 .31 .67 .31 .67 .31
Next-preferred $k$: medicine at U.Concepcion
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 76 23 255 182 255 182 255 372
Share listing <<8 choices .57 .47 .87 .96 .87 .96 .87 .96
$\underline{\delta}_{j,k}$ .71 .87 .21 .11 .21 .11 .21 .05
Next-preferred $k$: medicine at U.Santiago
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 61 14 123 38 123 38 123 352
Share listing <<8 choices .7 .07 .85 .97 .85 .97 .85 .96
$\underline{\delta}_{j,k}$ .86 1 .43 .51 .43 .51 .43 .05
Next-preferred $k$: sciences at PUC
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 70 32 72 58 77 147 77 400
Share listing <<8 choices .89 .18 .9 .94 .9 .96 .9 .93
$\underline{\delta}_{j,k}$ 1 .81 .97 .45 .9 .17 .9 .06
Next-preferred $k$: engineering at PUC
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 72 41 93 92 967 175 967 395
Share listing <<8 choices .51 .31 .86 .92 .96 .93 .96 .92
$\underline{\delta}_{j,k}$ .58 .38 .45 .17 .04 .08 .04 .03